OpenStudy (mendicant_bias):

Still chunking through Laplace-related problems. Posted below in a moment.

OpenStudy (mendicant_bias):

\[\mathcal{L}\left\{\vphantom{} \frac{e^{-2s}}{s^3}\right\}\]

OpenStudy (mendicant_bias):

My first thoughts looking at this are that a Heaviside function is going to be necessary in the end result, and the shift in that Heaviside is going to have to be equal to -2; \[a=-2.\]

OpenStudy (mendicant_bias):

This is the first problem where I've had to move backwards, in the other direction, when dealing with Heavisides in Laplace, so I'm a little bewildered by it; The next step I'm guessing would be some algebraic manipulation to get the s^3 term in the denominator to be in the appropriate form to have the same argument as the Heaviside function.

OpenStudy (mendicant_bias):

Whoops, supposed to be a Laplace Inverse! NOT a Laplace!

OpenStudy (mendicant_bias):

\[\mathcal{L^{-1}}\left\{\vphantom{}\frac{e^{-2s}}{s^3}\right\}\]

OpenStudy (sidsiddhartha):

yes your approach is correct\[L[u(t-a)]=\frac{ e^{-as} }{ s}\]

OpenStudy (mendicant_bias):

Alright, so I need an additional s^2, maybe I could break those up for the sake of looking at them separately:

OpenStudy (mendicant_bias):

\[\mathcal{L^{-1}}\left\{\vphantom{}\frac{e^{-2s}}{s^3}\right\}=\mathcal{L^{-1}}\left\{\vphantom{}\frac{e^{-2s}}{s}\frac{1}{s^2}\right\}\]

OpenStudy (sidsiddhartha):

\[L^{-1}\frac{ e^{-2s} }{ s^3 }=\int\limits_{0}^{t}\int\limits_{0}^{t}u(t-2)\\\] does that look correct

OpenStudy (mendicant_bias):

The double integral is really odd to me, but I can see why that's the case. I just wouldn't be writing it out that way, but I think I understand.

OpenStudy (sidsiddhartha):

now we can take take u(t-2) out of the integral and then integrate

OpenStudy (mendicant_bias):

How can we take U out of t is involved in the integral?

OpenStudy (mendicant_bias):

We're integrating with respect to t, yes?

OpenStudy (mendicant_bias):

No variable of integration.

OpenStudy (sidsiddhartha):

because \[u(t-2)=1 ~~for~~t>2 \]

OpenStudy (mendicant_bias):

Okay, so should the lower limits of integration be changed to 2, then?

OpenStudy (sidsiddhartha):

0 to t-2 i think

OpenStudy (sidsiddhartha):

its too much lagging :(

OpenStudy (sidsiddhartha):

so u should get \[L^{-1}\frac{ e^{-2s} }{ s^3 }=\frac{ 1 }{ 2 }(t-2)^2.u(t-2)\]

OpenStudy (sidsiddhartha):

seems ok?

OpenStudy (sidsiddhartha):

have u done a problem involving integrating heaviside before?

OpenStudy (sidsiddhartha):

if not i'm writing a example to clear things

OpenStudy (sidsiddhartha):

suppose u have --\[\int\limits_{-\infty}^{\infty}u(\tau-2).u(t -\tau-4)d \tau\\step ~1:u(t-\tau)=1,for~(\tau-2)>0~or~\tau>2\\step 2::u(t-\tau-4)=1,for~t-\tau-4>0~or~\tau<t-4\\so ~the~\int\limits~becomes\\f(t)=\int\limits_{2}^{t-4}1.1 d \tau=(t-6)u(t-6)\] like this

OpenStudy (sidsiddhartha):

still confused?

OpenStudy (mendicant_bias):

I haven't done a problem integrating heaviside before, this is my first; I'm looking at the example now, thank you!

OpenStudy (mendicant_bias):

Isn't this t minus tau stuff convolution?

OpenStudy (sidsiddhartha):

yes the idea comes from here

OpenStudy (sidsiddhartha):

i can give u another method but i dont know y'll understand it or not , i'm attaching a photo-

OpenStudy (sidsiddhartha):

wolfram gives the same result http://www.wolframalpha.com/input/?i=inverse+laplace+%5Be%5E%7B-2s%7D%2Fs%5E3%5D

OpenStudy (mendicant_bias):

Hey siddhartha, I'm guessing the character to left-is that gamma or something? http://i.imgur.com/IVSZwaN.pngf

OpenStudy (mendicant_bias):

Oh, "r" for "ramp"?

OpenStudy (mendicant_bias):

I've never seen a ramp function before, either; my book altogether doesn't mention it. Thanks for putting so much effort into all of this stuff, man. It's very generous of you.

OpenStudy (mendicant_bias):

My final answer will be in the form:\[\mathcal{L^{-1}}\left\{\vphantom{}\frac{e^{-2s}}{s^3}\right\}=f(t-2)\mathcal{U}(t-2),\]Where \[f(t-2)\]is the inverse laplace transform of \[\frac{1}{s^3}.\]This makes sense to me.

OpenStudy (mendicant_bias):

Correction, \[\mathcal{L^{-1}}\left\{\vphantom{}\frac{1}{s^3}\right\}=f(t),\]\[\mathcal{L^{-1}}\left\{\vphantom{}\frac{1}{s^3}\right\} \neq f(t-2).\]

OpenStudy (mendicant_bias):

\[\mathcal{L^{-1}}\left\{\vphantom{}\frac{1}{s^{3}}\right\}=\frac{1}{2}t^2\]

OpenStudy (mendicant_bias):

Accounting for the shift and multiplying by the unit step function, \[\frac{1}{2}(t-2)^2U(t-2)\] This is the correct answer, but I took a wholly different approach than yours; is my approach consistently valid, or am I making my own patterns that in fact don't exist?